Even though most corporate bonds in the United States make coupon payments semiannually, bonds issued elsewhere often have annual coupon payments. Suppose a German company issues a bond with a par value of €1,000, 9 years to maturity, and a coupon rate of 8 percent paid annually. If the yield to maturity is 9 percent, what is the current price of the bond?

  • €940.05

  • €1,066.47

  • $893.05

  • €1,030.00
Answer #1

Annual coupon=1000*8%=80

Hence current value=annual coupon*present value of annuity factor (9%,9) +1000*present value of discounting factor (9%,9)

=80*5.99524689+1000*0.46042778

=940.05(approx)

Note:

1.Present value of annuity=80[1-(1.09)^-9]/0.09

=80*5.99524689

2.present value of discounting factor=1000/1.09^9

=1000*0.46042778